OpenStudy (anonymous):

Find the volume of a pyramid with a square base with dimensions 100 mm by 100 mm and height of 75 mm. Enter the numerical answer only.

OpenStudy (anonymous):

I got 250,000 as an answer but I don't think it is right.

jimthompson5910 (jim_thompson5910):

250,000 is correct

OpenStudy (anonymous):

what about this question: Find the surface area of a cone with radius 30 in and slant height 50 in. Use 3.14 for pi. Enter the numerical answer only.

jimthompson5910 (jim_thompson5910):

you find the area of the base 100*100 = 10,000 then you multiply this with the height 10,000*75 = 750,000 finally, you divide this by 3 750,000/3 = 250,000

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

i cant figure it out... i have like 4 problems i have no clue on what to do...

OpenStudy (anonymous):

is the surface area 141371.67 and the volume 47123.89?

jimthompson5910 (jim_thompson5910):

surface area of cone SA = pi*r^2 + pi*r*s r = radius s = slant height

OpenStudy (anonymous):

SA=pi*30^2 + pi*30*50 SA=7539.82

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

if you used 3.14 for pi, you would get 7,536

jimthompson5910 (jim_thompson5910):

so 7,536 is your final answer

OpenStudy (anonymous):

Then how would you solve for the volume? V=1/3 (area of base) *(height)???

jimthompson5910 (jim_thompson5910):

good, you have the correct formula

jimthompson5910 (jim_thompson5910):

keep in mind that you need to find the height first

jimthompson5910 (jim_thompson5910):

it's not the same as the slant height btw

OpenStudy (anonymous):

so how would I do that?

jimthompson5910 (jim_thompson5910):

hint:  h = height r = radius s = slant height

jimthompson5910 (jim_thompson5910):

another hint h^2 + r^2 = s^2

OpenStudy (anonymous):

so h would be 40?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so the volume would be 100480?

jimthompson5910 (jim_thompson5910):

what is the area of the base

OpenStudy (anonymous):

7536?

jimthompson5910 (jim_thompson5910):

that was the overall surface area

OpenStudy (anonymous):

ohhhh.....

jimthompson5910 (jim_thompson5910):

the base is a circle with a radius of 30

OpenStudy (anonymous):

so how would I go about solving this... because the formula for area is l*w

jimthompson5910 (jim_thompson5910):

area of circle A = pi*r^2 r = 30 in this case

OpenStudy (anonymous):

so A=pi*30^2 A=2827.43 THEN.... V=1/3(area of base)(height) V=1/3(2827.43)(40) V=37699.11184???

jimthompson5910 (jim_thompson5910):

remember you're using 3.14 for pi, so pi*30^2 = 3.14*30^2 = 2826

jimthompson5910 (jim_thompson5910):

so the area of the base is 2826

OpenStudy (anonymous):

ohh i am using my pi button on my calculator

OpenStudy (anonymous):

I only have 2 other problems that I am really confused on in this quiz: 1: What is the volume of a pyramid that has a base with an area of 81 square feet and a height of 4 feet? Enter the numerical answer only. 2: Find the radius of the base of a cone with volume equal to nr010-1.jpg in³ and height equal to 12 in. Use 3.14 for pi. Enter the numerical answer only.

jimthompson5910 (jim_thompson5910):

that's ok, i was too until I realized they put in "Use 3.14 for pi."

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

What is the volume of a pyramid that has a base with an area of 81 square feet and a height of 4 feet? Enter the numerical answer only.

jimthompson5910 (jim_thompson5910):

Volume of Pyramid = (Area of base)*(height)/3

OpenStudy (anonymous):

so i would do 81*4/3

OpenStudy (anonymous):

which is 108?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Find the radius of the base of a cone with volume equal to nr010-1.jpg in³ and height equal to 12 in. Use 3.14 for pi. Enter the numerical answer only.

jimthompson5910 (jim_thompson5910):

what is nr010-1.jpg in³ all about??

OpenStudy (anonymous):

144pi

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

Volume of Cone = (area of base)*(height)/3 ... note: a cone is a special type of pyramid

jimthompson5910 (jim_thompson5910):

Volume of Cone = (area of base)*(height)/3 Volume of Cone = (pi*r^2)*(height)/3 V = (pi*r^2)*(h)/3 144pi = (pi*r^2)*(12)/3 solve for r

jimthompson5910 (jim_thompson5910):

hint: the pi terms will divide and cancel, so no need to worry about that

OpenStudy (anonymous):

so i will have 144/12*3?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

and then square root

jimthompson5910 (jim_thompson5910):

144pi = (pi*r^2)*(12)/3 144 = (12r^2)/3 .. .. .. r = ??

jimthompson5910 (jim_thompson5910):

oh wait, misread what you typed, yes you're on the right track

OpenStudy (anonymous):

144*3

OpenStudy (anonymous):

432/12

OpenStudy (anonymous):

square root of 36

OpenStudy (anonymous):

r =6

jimthompson5910 (jim_thompson5910):

very good